# Phase Transition for Budgeted Multi-Agent Synergy

### v1

@@ 3.3 Saturation floors and mixing depth
depth LLMSE $v_{L}$$v^{\star}(b,m,\rho)$$v^{\star}(b,m^{\prime},\rho)$shorter messageslonger messagesexponential shrink until saturation
Figure 3: Conceptual behavior of the tree recursion (9). Increasing depth reduces error rapidly at first (by repeatedly averaging), but improvement saturates at a floor $v^{\star}$ determined by communication distortion and shared correlation. Longer messages (larger $m$) reduce $\sigma_{c}^{2}(m)$, lowering the floor, but also reduce feasible fan-in due to the context constraint $bm\leq W$.
@@ 3.4 Scale-out versus scale-up under a fixed budget
$v_{L}\;\approx\;c\,x^{-\beta}\left(\frac{B}{x+c_{0}}\right)^{-t}=c\,B^{-t}\,x^{-\beta}(x+c_{0})^{t}.$
@@ 3.5 Design implications and a bridge to the binary analysis
The binary case is qualitatively different. Averaging is linear, so errors evolve smoothly toward an explicit floor. Majority vote discards magnitude and retains only a sign; when the incoming signal is too weak relative to noise and correlation, repeated majority aggregation can wash out information rather than denoise it. Section 4 formalizes this amplification–collapse transition and shows how the same measurable quantities ($\rho$, message length $m$, fan-in $b$, and budget allocation) determine whether organization helps or hurts on success/failure tasks.
@@ 4 Binary Core: Majority Aggregation and Phase Transitions
Binary success/failure tasks behave differently from the continuous warm-up because majority vote discards magnitude information and retains only a sign. That nonlinearity can either amplify weak signal across layers or wash it out entirely. As a result, hierarchical organization exhibits a sharp tipping point: below a threshold, deeper trees collapse to chance; above it, depth improves accuracy until it saturates.
Real multi-agent pipelines often mix graded intermediate signals (scores, confidences, rankings) with hard decisions (accept/reject, pass/fail). We therefore view the continuous and binary analyses as two canonical primitives that appear within richer protocols; the same measurable constraints (communication fidelity, shared failures, and context-limited fan-in) govern both.
@@ 4.1 Majority maps and correlated aggregation
From leaf votes to bias.
[…]
$\mu:=\mathbb{E}[\widehat{Y}Y],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$
In a tree, each internal node receives $b$ child votes and outputs an aggregated vote. To analyze such systems, we need to understand how bias transforms through one aggregation step.
@@
where ties are impossible because $b$ is odd. The induced majority map is the bias at the output:
@@
input bias uuoutput bias $f_{b}(u)$f=uf=ub=3b=3b=5b=5b=7b=7b=11b=11
Slope at origin: $f_{b}^{\prime}(0)=\frac{b\binom{b-1}{(b-1)/2}}{2^{b-1}}$    (grows as $O(\sqrt{b})$)
Figure 4: The majority aggregation map $f_{b}(u)$ for odd fan-in $b\in\{3,5,7,11\}$. Each curve shows how input bias $u$ transforms into output bias after majority voting. All curves pass through $(0,0)$ and $(1,1)$, with slope $f_{b}^{\prime}(0)>1$ at the origin—majority amplifies weak signals. Larger fan-in produces stronger amplification, but the gain grows only as $O(\sqrt{b})$.
@@ 4.2 Phase transition in deep trees: $\alpha_{\rho}\gtrless 1$
We assume odd fan-in $b$ to avoid ties in majority voting; even $b$ can be handled by specifying a tie-breaking rule (e.g., random) with only minor notational changes. Figure 6 illustrates the two regimes: when $\alpha_{\rho}<1$ the recursion map lies below the diagonal and $\mu_{t}$ collapses to 0, while when $\alpha_{\rho}>1$ the map crosses the diagonal at a stable fixed point $\mu^{\star}$, yielding amplification followed by saturation.
@@
$\mu_{t}$$T(\mu_{t})$$T(\mu)=\mu$$\mu^{\star}$$\alpha_{\rho}<1$$\alpha_{\rho}>1$
(a) Recursion map (schematic)
depth $t\mu_{t}$$\mu^{\star}$collapseamplify then saturate
(b) Trajectories (schematic)
Figure 6: A phase transition in deep majority-vote trees. Left: the recursion map $T(\mu)=f_{b,\rho}(\gamma\mu)$ compared to the diagonal. If $\alpha_{\rho}=T^{\prime}(0)<1$, the map stays below the diagonal and bias decays to 0. If $\alpha_{\rho}>1$, the map is initially above the diagonal, creating amplification, but boundedness and communication loss yield a stable fixed point $\mu^{\star}$. Right: corresponding trajectories from the same initialization.
@@ 4.3 Small-signal amplification and the organization exponent $s$
Fix $b,\rho,\gamma$ and let $T(u)=f_{b,\rho}(\gamma u)$ with $\alpha_{\rho}=T^{\prime}(0)$. For any $\eta\in(0,1)$, there exists $\delta=\delta(\eta;b,\rho,\gamma)\in(0,1]$ such that for all $u\in[0,\delta]$,
@@
$\mu_{L}\approx\mu_{0}\,\alpha_{\rho}^{L}=\mu_{0}\,N^{s},\qquad s:=\frac{\log\alpha_{\rho}}{\log b}.$
@@
This exponent is the main interface between organization and budget. Later, in Section 5, we compare $s$ to the single-agent scaling exponent $\beta$ to decide whether scale-out can beat scale-up under a fixed budget. At this stage, the key point is conceptual: $s$ increases with better communication ($\gamma(m)$), decreases with shared failures ($\rho$), and is limited by fan-in through $f_{b}^{\prime}(0)$.
@@ 4.4 Saturation and finite-depth guarantees: $L_{\mathrm{mix}}$ and clipped objectives
There are two phases. First, while $\mu_{t}$ is small, the map grows approximately linearly with factor $\alpha_{\rho}>1$, so reaching a constant-size neighborhood of $\mu^{\star}$ takes $O(\log(1/\mu_{0}))$ steps. Second, concavity implies $T^{\prime}(\mu)\leq\kappa<1$ on a neighborhood below $\mu^{\star}$, yielding a contraction argument that reaches relative error $\varepsilon$ in $O(\log(1/\varepsilon))$ additional steps. An explicit construction and constants appear in Appendix E. $\square$
@@
Section 5 uses the organization exponent $s$, the threshold $\alpha_{\rho}>1$, and the clipped objective (27) to derive topology and budget phase diagrams under context limits. Section 6 turns these results into a theory-guided design algorithm that outputs monotone communication design curves and compute allocations under a fixed budget.
5 Topology and Budget Phase Diagrams
Sections 34 characterized how information transforms through a single aggregation layer. This section turns those layer-wise results into system-level guidance: when centralized aggregation (star) is feasible, when hierarchy (tree) is necessary, and how these choices interact with a fixed budget $B$ under a context window constraint $W$. The resulting phase diagrams organize a large design space into regimes governed by the environment parameters of Section 2 (the single-agent scaling exponent $\beta$, shared-failure correlation $\rho$, and communication fidelity $\gamma(m)$) together with the context constraint $W$.
We focus on the binary setting and use the recursion and organization exponent developed in Section 4; when helpful, we connect back to the continuous warm-up to build intuition.
5.1 Star as an upper bound without constraints
@@
5.2 Why chains degrade without new evidence
@@
Absent new evidence, chains are not a mechanism for synergy. If a task is inherently sequential, chains may still be necessary, but then one must model the arrival of new information explicitly. If intermediate agents inject new information or perform error-correcting verification, the effective bias can increase; Proposition 8 isolates the relay-only communication effect. Our phase diagrams therefore focus on star versus tree, which represent the two competing strategies for aggregating many leaf signals under context constraints.
5.3 Context-limited star versus hierarchical trees
@@
[…]
Figure 7: Schematic performance comparison under a fixed context window $W$. Star (orange) aggregates NN leaves but saturates when NN hits the ceiling $N_{\max}=\lfloor W/m\rfloor$. Tree (blue) distributes the constraint locally ($bm\leq W$), enabling $N=b^{L}$ leaves. When $s>\beta$, trees overtake at budget $B_{\mathrm{crit}}$ and continue with slope $s$ until reaching $\mu^{\star}$.
@@
[…]
Figure 8: A schematic binary phase diagram for hierarchical trees at fixed fan-in $b$ and single-agent scaling exponent $\beta$. The lower boundary $\alpha_{\rho}=1$ (Theorem 4) separates regimes where deep hierarchies collapse to chance from regimes where they amplify accuracy. Within the amplifying regime, the stronger condition $\alpha_{\rho}>b^{\beta}$ (equivalently $s>\beta$) marks where scaling out can beat scaling up in the budgeted growth regime. Both boundaries move upward as $\rho$ increases. Since $\gamma(m)\leq 1$, the $\alpha_{\rho}=b^{\beta}$ boundary may exit the feasible range at large $\rho$; the schematic clips it at $\gamma=1$, truncating the synergy region.
@@
5.4 Budget thresholds for synergy: conditions on $s$ and $\beta$
We now connect topology to budget. The question is not whether a deep tree can amplify weak votes in isolation (Section 4), but whether it can do so more efficiently than simply making a single agent stronger. This is where the comparison between the organization exponent $s$ and the single-agent scaling exponent $\beta$ becomes decisive.
@@ 5.4 Budget thresholds for synergy: conditions on $s$ and $\beta$
We work in the small-signal regime where the single-agent scaling law is well approximated by
$\mu_{0}(x)=g(x)\approx kx^{\beta},$
and where the tree has not yet saturated (so Theorem 5 applies). For a full $b$-ary tree of depth $L$ with $N=b^{L}$ leaves, the small-signal prediction gives
$\mu_{L}\approx\mu_{0}(x)\,\alpha_{\rho}^{L}=\mu_{0}(x)\,N^{s},\qquad s=\frac{\log\alpha_{\rho}}{\log b}.$ (32)
Under the budget model of Section 2.5, the total cost is approximately
$B\approx N\big(x+c_{0}(b,m)\big),\qquad c_{0}(b,m)=2m\frac{b}{b-1},$
so $N\approx B/(x+c_{0})$. Substituting into (32) yields the design objective
$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s}.$ (33)
This objective is interpretable: $x^{\beta}$ captures scale-up, and $(B/(x+c_{0}))^{s}$ captures scale-out through the number of leaves.
@@
Fix $(b,m,\rho)$ and suppose $s(b,m,\rho)>\beta$. Then the growth objective (33) is maximized at
$x^{\star}\;=\;\frac{\beta}{s-\beta}\,c_{0}(b,m),$ (34)
@@
$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa(s,\beta)\,c_{0}(b,m)^{\beta-s}\,B^{s},\qquad\kappa(s,\beta):=\beta^{\beta}(s-\beta)^{\,s-\beta}s^{-s}.$ (35)
If $s\leq\beta$, the objective (33) is maximized by pushing $x$ as large as allowed (scale up rather than scale out).
@@
For $s>\beta$, maximize $\log\mu_{\mathrm{grow}}=\beta\log x+s\log B-s\log(x+c_{0})$ over $x>0$. Setting the derivative to zero gives (34). Substituting back yields (35). If $s\leq\beta$, the derivative is nonnegative for all $x$, so the objective increases with $x$. Details are routine and given in Appendix D. $\square$
@@
A single-agent baseline that spends the full budget on compute achieves (in the same small-signal regime)
$\mu_{\mathrm{single}}(B)\approx k\,B^{\beta}.$ (36)
Comparing (35) and (36) yields a simple conclusion: in the growth regime, a hierarchy can only outpace the single agent when $s>\beta$. Moreover, when $s>\beta$, there is an explicit budget threshold beyond which the optimized tree dominates the single-agent scaling curve.
@@
Fix $(b,m,\rho)$ with $s>\beta$, and define $\kappa(s,\beta)$ as in (35). Ignoring saturation, a sufficient condition for the optimized tree prediction $\mu_{\mathrm{grow}}^{\star}(B;b,m)$ to exceed the single-agent prediction $\mu_{\mathrm{single}}(B)$ is
$B\;\geq\;B_{\mathrm{crit}}(b,m)\;:=\;c_{0}(b,m)\,\kappa(s,\beta)^{-\frac{1}{s-\beta}}.$ (37)
@@
Equation (37) is a concrete, testable statement. It predicts that scale-out synergy is harder when coordination is expensive (large $c_{0}$), when the organization exponent barely beats the scaling exponent (small $s-\beta$), or when the system is close to the amplification boundary (small $s$). In Section 4.4 we introduced clipped objectives to account for saturation; when saturation matters, the threshold still indicates whether there exists a budget range in which scale-out should help before the fixed point dominates.
[…]
Figure 9: A schematic budget phase diagram under a fixed context window $W$. A star cannot incorporate more than $N_{\max}(m)$ leaves, so beyond a point its gains come primarily from making leaves stronger (scale-up), leading to a $B^{\beta}$-type growth until task-level saturation. A supercritical tree with organization exponent $s>\beta$ can exhibit a steeper $B^{s}$ growth regime before saturating at $\mu^{\star}$. The budget threshold $B_{\mathrm{crit}}$ marks where optimized scale-out can overtake single-agent scaling in the growth regime.
5.5 A universal upper bound on amplification exponents

### v2

@@ 3.3 Saturation floors and mixing depth
Figure 4: Continuous averaging recursion through a $b$-ary hierarchy with equal-correlation $\rho$. The MSE $v_{L}$ contracts approximately exponentially with depth $L$ until it reaches the communication-limited floor $v^{\star}(b,m,\rho)=\sigma_{c}^{2}(m)/\bigl((b-1)(1-\rho)\bigr)$. Longer messages reduce $\sigma_{c}^{2}(m)$ and therefore lower the attainable floor.
@@ 3.4 Scale-out versus scale-up under a fixed budget
$v_{L}\;\approx\;c\,x^{-\beta}\left(\frac{B}{x+c_{0}}\right)^{-t}=c\,B^{-t}\,x^{-\beta}(x+c_{0})^{t}.$ (22)
@@ 3.5 Design implications and a bridge to the binary analysis
The binary case is qualitatively different. Because majority vote is nonlinear, it can fundamentally change how errors propagate under repeated aggregation. Section 4 analyzes this effect and shows that hierarchical organization exhibits a sharp amplification–collapse transition on success/failure tasks.
@@ 4 Binary Core: Majority Aggregation and Phase Transitions
We begin by characterizing how one majority layer transforms weak bias, then incorporate correlated errors and lossy communication to obtain a phase transition and budget-relevant exponents for deep trees.
@@ 4.1 Majority maps and correlated aggregation
The central object in the binary analysis is the one-step aggregation map that describes how an internal node transforms the biases of its child votes into an output bias.
@@
where ties are impossible because $b$ is odd; even $b$ can be handled by specifying a tie-breaking rule (e.g., random) with only minor notational changes. For the same reason, we will assume an odd $b$ in our future sections. The induced majority map is the bias at the output:
@@
Figure 5: Majority amplification map $f_{b}(u)$ for odd fan-in $b$ (shown for $b\in\{3,5,7,11\}$), where $u$ is the input bias and $f_{b}(u)=2\Pr(\mathrm{Bin}(b,(1+u)/2)\geq(b+1)/2)-1$. The diagonal $f(u)=u$ is shown for reference. The slope at the origin, $f_{b}^{\prime}(0)=\frac{b\binom{b-1}{(b-1)/2}}{2^{\,b-1}}$, grows on the order of $\sqrt{b}$, capturing how majority increasingly amplifies weak signals as fan-in grows.
@@ 4.2 Phase transition in deep trees: $\alpha_{\rho}\gtrless 1$
Figure 7 illustrates the two regimes: when $\alpha_{\rho}<1$ the recursion map lies below the diagonal and $\mu_{t}$ collapses to 0, while when $\alpha_{\rho}>1$ the map crosses the diagonal at a stable fixed point $\mu^{\star}$, yielding amplification followed by saturation.
@@
(a) Recursion map $T(\mu)$ vs. the diagonal.
(b) Iterates $\mu_{t+1}=T(\mu_{t})$ from the same initialization.
Figure 7: Binary phase transition under the recursion map $T(\mu)=f_{b,\rho}(\gamma\mu)$ where $f_{b,\rho}(u)=\rho u+(1-\rho)f_{b}(u)$. In the subcritical regime ($\alpha_{\rho}<1$) trajectories collapse to $\mu=0$. In the supercritical regime ($\alpha_{\rho}>1$) a stable fixed point $\mu^{\star}>0$ emerges and trajectories amplify then saturate at $\mu^{\star}$.
@@ 4.3 Small-signal amplification and the organization exponent $s$
Fix odd $b\geq 3$, $\rho\in[0,1)$, and channel reliability $\gamma=\gamma(m)\in(0,1]$. Let $T(u)=f_{b,\rho}(\gamma u)$ and $\alpha_{\rho}:=T^{\prime}(0)$. For any $\eta\in(0,1)$, there exists $\delta=\delta(\eta;b,\rho,\gamma)\in(0,1]$ such that for all $u\in[0,\delta]$,
@@
$\mu_{L}\approx\mu_{0}\,\alpha_{\rho}^{L}=\mu_{0}\,N^{s},\qquad s:=\frac{\log\alpha_{\rho}}{\log b}.$ (34)
@@
This exponent is the main interface between organization and budget. Later, in Section 4.5, we compare $s$ to the single-agent scaling exponent $\beta$ to decide whether scale-out can beat scale-up under a fixed budget. At this stage, the key point is conceptual: $s$ increases with better communication ($\gamma(m)$), decreases with shared failures ($\rho$), and is limited by fan-in through $f_{b}^{\prime}(0)$.
@@ 4.4 Saturation and finite-depth guarantees: $L_{\mathrm{mix}}$ and clipped objectives
Write the recursion as $\mu_{t+1}=T(\mu_{t})$ with $T(u)=f_{b,\rho}(\gamma u)$ and a stable fixed point $\mu^{\star}$ when $\alpha_{\rho}=T^{\prime}(0)>1$. There are two phases. First, while $\mu_{t}$ remains in the small-signal band, Theorem 5 gives $(1-\eta)\alpha_{\rho}^{t}\mu_{0}\leq\mu_{t}\leq\alpha_{\rho}^{t}\mu_{0}$, so reaching a constant-sized neighborhood of $\mu^{\star}$ takes $O(\log(1/\mu_{0}))$ steps. Second, stability implies $T^{\prime}(\mu^{\star})<1$; by continuity there exists a neighborhood below $\mu^{\star}$ on which $T^{\prime}(\mu)\leq\kappa<1$, so the recursion is a contraction and the error shrinks geometrically, reaching relative error $\varepsilon$ in $O(\log(1/\varepsilon))$ additional steps. Appendix E provides an explicit construction and constants. $\square$
@@
Section 4.5 uses the organization exponent $s$, the threshold $\alpha_{\rho}>1$, and the clipped objective (35) to derive topology and budget phase diagrams under context limits. Section 5 turns these results into a theory-guided design algorithm that outputs monotone communication design curves and compute allocations under a fixed budget.
4.5 Topology and Budget Phase Diagrams
Building on the binary recursion and organization exponent from Section 4, we translate the layer-wise results into system-level guidance under a fixed budget $B$ and context window $W$. We ask when a context-limited star is feasible, when hierarchy is necessary, and when scaling out can beat scaling up before saturation. The resulting regimes can be summarized as phase diagrams in the measurable environment parameters $\beta$, $\rho$, and $\gamma(m)$ (Section 2), together with the context constraint $W$.
4.5.1 Star as an upper bound without constraints
@@
4.5.2 Why chains degrade without new evidence
@@
A relay-only chain strictly attenuates signal (Proposition 8); it helps only when intermediate steps add new evidence (tools, retrieval, verification) rather than passing along the same information. We therefore focus on star versus tree as the two scalable evidence-aggregation strategies under context limits.
4.5.3 Context-limited star versus hierarchical trees
@@
Figure 8: Star vs. tree vs. single-agent performance under a fixed context window $W$. A star can increase fan-in NN until it hits the context ceiling $N_{\max}=\lfloor W/m\rfloor$ (defining $B_{\mathrm{sat}}$), after which further budget can only improve per-leaf compute (scaling like $B^{\beta}$). A hierarchical tree can continue increasing the effective number of leaves by adding depth, exhibiting a growth exponent $s=\log(\alpha_{\rho})/\log b$ in the growth regime; when $s>\beta$, trees eventually dominate beyond $B_{\mathrm{crit}}$.
@@
Figure 9: Phase regions in $(\rho,\gamma)$ for fixed fan-in $b$ and single-agent scaling exponent $\beta$. The boundary $\alpha_{\rho}=1$ separates collapse ($\alpha_{\rho}\leq 1$) from amplification. The boundary $\alpha_{\rho}=b^{\beta}$ separates mere amplification (no synergy, $s\leq\beta$) from synergy ($s>\beta$), where hierarchical scale-out can outpace single-agent scaling. Design levers such as increasing message length $m$ (raising $\gamma$) or reducing shared-failure correlation $\rho$ move systems toward the synergy region.
@@
4.5.4 Budget thresholds for synergy: conditions on $s$ and $\beta$
The layer-wise analysis of Section 4 yields an organization exponent $s$ that governs how signal grows with depth before saturation. Here we translate that picture into a fixed-budget comparison: under the same total budget $B$, when can scaling out with a tree outperform scaling up a single agent?
@@ 4.5.4 Budget thresholds for synergy: conditions on $s$ and $\beta$
Assume the per-leaf scaling curve admits a local exponent $\beta$ in the operating range (Section 2.2), so that $\mu_{0}(x)=g(x)\approx kx^{\beta}$ when leaf bias is small. In the amplification band of Theorem 5 (i.e., before the recursion approaches $\mu^{\star}$), a depth-$L$ full $b$-ary tree satisfies
$\mu_{L}\;\approx\;\mu_{0}(x)\,\alpha_{\rho}(b,m)^{L}\;=\;\mu_{0}(x)\,N^{s(b,m,\rho)},\qquad N=b^{L},$ (40)
where $\alpha_{\rho}(b,m)$ is defined in (29) and $s(b,m,\rho)=\log\alpha_{\rho}(b,m)/\log b$ as in (34). With the tree budget model $B\approx N\big(x+c_{0}(b,m)\big)$ from Section 2.5, we have $N\approx B/(x+c_{0})$, giving the growth surrogate
$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s(b,m,\rho)}.$ (41)
We use (41) only as a pre-saturation diagnostic; in Section 4.4 we clip predictions at the fixed point $\mu^{\star}(b,m)$.
@@
Fix $(b,m,\rho)$ and suppose $s(b,m,\rho)>\beta$. Then the objective (41) is maximized at
$x^{\star}(b,m)=\frac{\beta}{s(b,m,\rho)-\beta}\,c_{0}(b,m),$ (42)
@@
$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa\big(s(b,m,\rho),\beta\big)\,c_{0}(b,m)^{\beta-s(b,m,\rho)}\,B^{s(b,m,\rho)},$ (43)
where $\kappa(s,\beta)=\beta^{\beta}(s-\beta)^{s-\beta}s^{-s}$. If $s(b,m,\rho)\leq\beta$, then (41) is (weakly) increasing in $x$, so the growth surrogate favors scale up over scale out.
@@
The optimization is one-dimensional: maximize $\log\mu_{\mathrm{grow}}=\beta\log x-s\log(x+c_{0})+s\log B$. Setting the derivative to zero yields (42), and substituting gives (43). This is the same calculus as in the continuous warm-up (Section 3.4), with the continuous exponent $t$ replaced by the binary exponent $s$; Appendix D.2 gives the full derivation. $\square$
@@
Spending the whole budget on a single agent yields
$\mu_{\mathrm{single}}(B)\;=\;\mu_{0}(B)\;\approx\;k\,B^{\beta}.$ (44)
Comparing (43) with (44) shows that exponent-level scale-out synergy in the growth regime requires $s>\beta$ (equivalently $\alpha_{\rho}>b^{\beta}$).
@@
Fix $(b,m,\rho)$ with $s(b,m,\rho)>\beta$. If
$B\;\geq\;B_{\mathrm{crit}}(b,m)\;\approx\;c_{0}(b,m)\,\kappa\big(s(b,m,\rho),\beta\big)^{-1/(s(b,m,\rho)-\beta)},$ (45)
then the optimized growth prediction (43) exceeds the single-agent scaling ($4$).
@@
Equation (45) makes the trade-off explicit: synergy is harder when coordination is expensive (large $c_{0}$) or when $s$ only barely exceeds $\beta$. Saturation can truncate the growth regime, but the threshold still indicates whether a budget window exists in which scale-out should help before the fixed point $\mu^{\star}$ dominates.
4.5.5 A universal upper bound on amplification exponents